### v2

@@ Abstract
Existing learning-based stereo image codec adopt sophisticated transformation with simple entropy models derived from single image codecs to encode latent representations. However, those entropy models struggle to effectively capture the spatial-disparity characteristics inherent in stereo images, which leads to suboptimal rate-distortion results. In this paper, we propose a stereo image compression framework, named CAMSIC. CAMSIC independently transforms each image to latent representation and employs a powerful decoder-free Transformer entropy model to capture both spatial and disparity dependencies, by introducing a novel content-aware masked image modeling (MIM) technique. Our content-aware MIM facilitates efficient bidirectional interaction between prior information and estimated tokens, which naturally obviates the need for an extra Transformer decoder. Experiments show that our stereo image codec achieves state-of-the-art rate-distortion performance on two stereo image datasets Cityscapes and InStereo2K with fast encoding and decoding speed.
Keywords:
Stereo Image Compression Masked Image Modeling Transformer Entropy Model
footnotetext: ${}^{\ast}$his work was partially done when Xinjie Zhang interned at SenseTime Research.footnotetext: ${}^{\dagger}$orresponding Author.
1 Introduction
@@
Stereo Image Codec (SIC) compresses a pair of stereoscopic images captured from distinct viewpoints by the same camera. SIC has attracted significant interest due to the growing demand for high-quality stereo image transmission and storage across applications such as autonomous driving [48], virtual reality streaming [20] and video surveillance [37]. By leveraging the inherent correlation between views, they achieve higher coding efficiency compared with single image compression.
Over the past decades, classical multi-view image coding standards, such as H.264-based MVC [41] and H.265-based MV-HEVC [39], have catalyzed the development of learning-based stereo image compression approaches [29, 19, 45, 49, 18]. Given a reference view, these methods, rooted in the predictive coding paradigm, employ disparity-compensated prediction in either pixel or feature space to compress current view. Additionally, recent advancements have introduced a bidirectional coding framework [27, 44] that utilizes a cross-attention mechanism, further enhancing content correlation exploitation between stereo images. Albeit remarkable, the odyssey of stereo image compression research concentrates on producing compact representations through various information flows or network designs before entropy coding. However, an orthogonal direction, which reduces redundancy by devising a superior spatial-disparity entropy model for entropy coding, is rarely explored.
Previous learning-based SIC approaches often adapt existing single image compression solutions, such as hyperprior [6] and auto-regressive prior [33], into conditional formats. These adapted entropy models, typically based on convolutional neural network (CNN) architectures, aim to capture spatial-disparity correlations. However, CNNs are inherently limited by their local receptive fields, which hampers their ability to model long-range dependencies. This limitation often leads to a rough estimation of probability distribution and inferior compression performance, especially when the disparity distance between stereo images is large. Consequently, we argue that better coding gains can be achieved as long as we find a more effective spatial-disparity entropy model.
@@
Figure 1: Comparison of two different mask image modeling Transformer entropy models. Tokens within the same purple rectangle have bidirectional interactions. The vanilla content-irrelevant masked image modeling uses uniform [MASK] tokens to fill the unestimated positions and relies on an inefficient encoder-decoder Transformer architecture to propagate the prior information. Differently, our content-aware masked image modeling replaces the uninformative [MASK] tokens with proposed content-aware tokens, which allows us to use an efficient decoder-free architecture with more sufficient information interactions.
@@
In this paper, we present a stereo image compression framework that centers on a powerful Transformer entropy model to leverage spatial-disparity dependencies between current and previously decoded views. This framework incorporates a straightforward image encoder-decoder pair from [23] to extract latent representations for each view. As illustrated in Fig. 1 (a), we introduce the vanilla masked image modeling (MIM) style [13] to construct a content-irrelevant encoder-decoder Transformer entropy model. This model captures spatial information from the context of decoded tokens and disparity information from previous decoded view. Since the bi-directional self-attention in Transformer allows [MASK] tokens to use context from both directions, it supports a non-sequential auto-regressive decoding process and decodes all tokens in the images in a few steps, achieving a good performance-speed trade-off [13].
@@
Nevertheless, the vanilla MIM-based Transformer entropy model still exhibits two significant shortcomings. Firstly, it uses uniform [MASK] tokens to indicate the positions of the tokens whose probability distributions are yet to be estimated, which are irrelevant to the video content and uninformative to entropy coding. As a result, it wastes much computation and leads to inferior coding gains. Secondly, it needs an extra Transformer decoder to propagate the prior information via cross-attention. Such a complicated encoder-decoder Transformer architecture only permits an inefficient unidirectional propagation from the prior information to the input tokens of the Transformer decoder. To overcome these challenges, we develop a novel Content-Aware Masked image modeling Transformer entropy model for Stereo Image Compression (CAMSIC) shown in Fig. 1 (b). Specifically, we introduce content-aware tokens generated by prior information to replace the uninformative [MASK] tokens, forming a new content-aware masked image modeling style, which brings three major benefits. First, when updating the features of the estimated tokens, because interacting with the proposed content-aware tokens provides useful information about each specific position, it prevents the distraction of previous uninformative [MASK] tokens. Second, it allows a bidirectional information flow between prior information and estimated tokens via self-attention. Compared with static prior information in the vanilla MIM, content-aware tokens, carrying specific prior information, can also be iteratively updated. Thus, it can effectively capture the context information from already estimated tokens and reduce the estimation uncertainty of remaining tokens. Third, since our design naturally embeds prior information propagation into each self-attention operation, we discard the whole Transformer decoder and devise an encoder-only Transformer architecture with low computation overhead.
In summary, our main contributions are three-fold:
@@
    We introduce a learning-based stereo image compression framework with a simple image encoder-decoder pair, which focuses on using a elegantly neat but powerful Transformer entropy model based on masked image modeling to exploit the relationship between the left and right images.
@@
    Experimental results show that our proposed method with lower encoding and decoding latency significantly outperforms existing learning-based stereo image compression methods. When compared with the state-of-the-art method ECSIC [44], our method saves around 8.512% and 0.629% bitrates in terms of PSNR on Cityscapes and InStereo2K datasets, respectively. Detailed ablation studies and analyses validate the effectiveness of each proposed component.
2 Related Works
2.1 Single Image Compression
@@
Traditional standard image codecs, such as JPEG [42], JPEG2000 [36], BPG [9], and VVC-intra [12], typically employ a three-step process involving transformation, quantization, and entropy coding. Regrettably, these steps are optimized independently, which may compromise overall compression performance. In contrast, recent learning-based image compression [5, 6, 33, 16, 15, 51, 14, 52, 23, 31] have demonstrated remarkable success, outperforming traditional methods in terms of rate-distortion (RD) performance. These modern approaches utilize nonlinear transformations (e.g., generalized divisive normalization [4], attention mechanism [16, 15], and stacks of residual bottleneck blocks [23]) to generate concise latent representations. Moreover, a series of advanced entropy models, such as factorized prior [5], hyperprior [6], and auto-regressive context prior [33], are employed to accurately approximate the probability distribution of quantized latent representations. These existing works serve as foundational elements for our stereo image compression solution.
2.2 Stereo Image Compression
@@
Traditional multi-view image codecs, such as H.264-based MVC [41] and H.265-based MV-HEVC [39], typically compress stereo images using disparity compensation prediction, which enjoys high compression efficiency but heavily depends on prior knowledge to design hand-crafted modules. Recently, learning-based stereo image compression methods have shown substantial improvements in compression performance. These existing works can be roughly classified into two categories: (1) Unidirectional coding [29, 19, 45, 49, 18] begins with the prediction of a disparity-compensated view in either pixel or feature space, followed by the compression of residuals between the current and predicted views. (2) Bidirectional coding [27, 44] incorporates a cross-attention mechanism within both the encoder and decoder to exploit the mutual information across stereo images comprehensively. Apart from the above mentioned joint encoding methods, some recent works [50, 35] have ventured into distributed multi-view image coding by using cross-attention alignment to achieve competitive RD performance.
Contrary to the prevailing focus on intricate nonlinear transformations by exploring different network structures or information flows, we investigate the application of the Transformer architecture in entropy coding and propose a gracefully simple but powerful Transformer entropy model for stereo image compression. It effectively exploits spatial-disparity correlations, which eliminates the need for complex feature extraction or warping operations, thereby streamlining the compression process and establishing a neat yet potent coding framework.
2.3 Masked Image Modeling
@@
Inspired by the masked language modeling [25] in natural language processing, masked image modeling (MIM) has been recently applied to representation learning for vision tasks [24, 47, 30]. It has shown great success on various downstream tasks, including image generation [13, 28], action recognition [40, 21], video compression [46] and video prediction [22]. In this paper, we adapt MIM technique for stereo image compression. Note that the vanilla MIM style uses a uniform [MASK] token to take the places of the tokens whose probability distribution are yet estimated, which wastes much computation on the content-irrelevant [MASK] tokens, thus leading to sub-optimal coding gains. In addition, it resorts to an extra Transformer decoder to propagate the prior information in a unidirectional manner. To address these limitations in the vanilla MIM, we develop a novel content-aware MIM style, which replaces the uninformative [MASK] tokens with content-aware tokens. It not only enables a bidirectional interaction between the prior information and estimated tokens, but also allows us to design an efficient decoder-free Transformer entropy model.
@@
3 Method
3.1 Overview
Our stereo image compression framework CAMSIC is shown in Fig. 2. Let $\mathcal{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2}\}$ denotes a pair of stereo images, where $\boldsymbol{x}_{v}\in\mathbb{R}^{H\times W\times 3}$ is the image at view $v$ with height $H$ and width $W$. An image encoder $E$ is used to compress each individual view $\boldsymbol{x}_{v}$ to a latent representation $\boldsymbol{y}_{v}\in\mathbb{R}^{h\times w\times d}$, where $h=H/16$, $w=W/16$ and $d$ is the latent dimension. $\boldsymbol{y}_{v}$ is then quantized to $\boldsymbol{\hat{y}}_{v}$ by the quantizer $Q$ as [34]. Finally, we feed $\boldsymbol{\hat{y}}_{v}$ to an image decoder $D$ to reconstruct the view $\boldsymbol{\hat{x}}_{v}$. Formally, the overall coding procedure can be formulated as
@@ 3.1 Overview
where $\mathcal{U}(\cdot)$ is uniform noise and *** denotes convolution. When encoding the first view, we expand a learned embedding with the dimension as $\mathbb{R}^{1\times 1\times d}$ to get $\boldsymbol{\hat{y}}_{-1}\in\mathbb{R}^{h\times w\times d}$. In this paper, we refer to prior information as the combination of disparity prior $\boldsymbol{\hat{y}}_{v-1}$ and hyperprior $\boldsymbol{\hat{z}}_{v}$ to distinguish the spatial context from other priors. By leveraging complementary priors, the conditional entropy model can estimate more accurate probability distributions.
3.2 Content-irrelevant Encoder-decoder Transformer Entropy Model
@@
Encoder-decoder Transformer. Since these [MASK] tokens are irrelevant to the video content, a Transformer decoder is required to propagate the content-related prior information to the [MASK] tokens through cross-attention. Fig. 3 (a) shows the architecture of an encoder-decoder Transformer entropy model based on content-irrelevant MIM. Specifically, we use a prior fusion network and a Transformer encoder to integrate the hyperprior $\boldsymbol{\hat{z}}_{v}$ and disparity prior $\boldsymbol{\hat{y}}_{v-1}$ for producing prior features. During training, we randomly generate a Boolean mask to composite the spatial context $\boldsymbol{s}_{v}$. The prior features and unmasked spatial tokens convey their information to the [MASK] tokens via cross-attention and self-attention in the Transformer decoder, receptively. Finally, we feed the feature $\boldsymbol{f}_{v}^{m}$ produced by the Transformer decoder into an entropy parameter module to estimate the probability distribution parameters $(\boldsymbol{\mu}_{v}^{m},\boldsymbol{\sigma}_{v}^{m})$.
Figure 3: Structures of content-irrelevant encoder-decoder Transformer entropy model and our content-aware decoder-free Transformer entropy model. HPD indicates hyperprior decoder. EPM means entropy parameter module. $\boldsymbol{m}$ is a mask randomly generated during training. (a) By combining the [MASK] token and previously estimated tokens in $\boldsymbol{\hat{y}}_{v}$, a composite context $\boldsymbol{s}_{v}$ is synthesized, which is then sent to the Transformer decoder to estimate the Gaussian distribution parameters $(\boldsymbol{\mu}_{v}^{m},\boldsymbol{\sigma}_{v}^{m})$ for masked tokens. (b) The [MASK] token is replaced by the context-aware tokens $\boldsymbol{c}_{v}$ to formulate the composite context $\boldsymbol{u}_{v}$. Then we feed $\boldsymbol{u}_{v}$ into Transformer encoder to output the estimated Gaussian distribution parameters $(\boldsymbol{\mu}_{v}^{m},\boldsymbol{\sigma}_{v}^{m})$.
3.3 Content-aware Decoder-free Transformer Entropy Model
Content-aware Masked Image Modeling. The vanilla MIM uses a content-irrelevant [MASK] token to occupy the places of the tokens whose probability distribution are yet estimated, which is not the most efficient and effective choice for entropy coding. On the one hand, interacting with the uninformative [MASK] tokens may distract the specific features of the estimated tokens during the repetitive self-attention operations. On the other hand, the encoder-decoder Transformer architecture is complicated but only allows a unidirectional propagation from prior information to the input tokens of the Transformer decoder. The prior information is fixed once extracted and cannot utilize the updated spatial context information from the already estimated tokens. Consequently, it is difficult for the entropy model to take full advantage of the power of Transformer architecture.
@@ 3.3 Content-aware Decoder-free Transformer Entropy Model
Decoder-free Transformer. Since the prior information has sufficiently interacted with the estimated tokens in the Transformer encoder, an explicit Transformer decoder is no longer needed. Thus, we discard the whole Transformer decoder and build an encoder-only Transformer architecture. As shown in Fig. 3 (b), it comprises two parts, i.e., a prior generation network that produces content-aware tokens $\boldsymbol{c}_{v}$ based on different priors and an decoder-free Transformer that estimates the probability distribution of the tokens. It is worth noting that we enable relative position embedding [32] in the Transformer and assign two identity embeddings to the token sequence to further distinguish different types of tokens, which can effectively promote the token interactions.
3.4 Training and Inference
@@
In the inference phase, the probability distributions of the tokens are iteratively estimated following a sinusoidal decoded ratio scheduler as [13]. In each iteration, only a fraction of tokens with the smallest entropy is selected as the estimated tokens whose probability distributions are determined afterwards. More details about iterative decoding process are available in the appendix.
Quantization. During training, since the quantization operation is not differentiable, we leverage a mixed quantizer to allow end-to-end optimization. To be specific, it adds $\boldsymbol{y}_{v}$ and a uniform noise $\mathcal{U}(-\frac{1}{2},\frac{1}{2})$ together as the input to the entropy model, but rounds $\boldsymbol{y}_{v}$ as the input to the decoder based on straight-through estimation [10]. We refer readers to [34] for more details.

### v3

@@ Abstract
Existing learning-based stereo image codec adopt sophisticated transformation with simple entropy models derived from single image codecs to encode latent representations. However, those entropy models struggle to effectively capture the spatial-disparity characteristics inherent in stereo images, which leads to suboptimal rate-distortion results. In this paper, we propose a stereo image compression framework, named CAMSIC. CAMSIC independently transforms each image to latent representation and employs a powerful decoder-free Transformer entropy model to capture both spatial and disparity dependencies, by introducing a novel content-aware masked image modeling (MIM) technique. Our content-aware MIM facilitates efficient bidirectional interaction between prior information and estimated tokens, which naturally obviates the need for an extra Transformer decoder. Experiments show that our stereo image codec achieves state-of-the-art rate-distortion performance on two stereo image datasets Cityscapes and InStereo2K with fast encoding and decoding speed. Code is available at https://github.com/Xinjie-Q/CAMSIC.
Introduction
@@
Stereo Image Codec (SIC) compresses a pair of stereoscopic images captured from distinct viewpoints by the same camera. SIC has attracted significant interest due to the growing demand for high-quality stereo image transmission and storage across applications such as autonomous driving (Yin et al. 2020), virtual reality (Fehn 2004) and video surveillance (Stepanov and Tishchenko 2016). By leveraging the inherent correlation between views, they achieve higher coding efficiency compared with single image codecs.
@@
Figure 1: Comparison of two different mask image modeling Transformer entropy models. [MASK] token is a learnable parameter, which is irrelevant to specific image content. Tokens within the same red rectangle have bidirectional interactions. The vanilla content-irrelevant masked image modeling uses uniform [MASK] tokens to fill the unestimated positions and relies on an inefficient encoder-decoder Transformer to propagate the prior information. Differently, our content-aware masked image modeling replaces the uninformative [MASK] tokens with proposed content-aware tokens, which enables an efficient decoder-free Transformer architecture with more sufficient information interactions.
Over the past decades, classical multi-view image coding standards, such as H.264-based MVC (Vetro, Wiegand, and Sullivan 2011) and H.265-based MV-HEVC (Tech et al. 2015), have catalyzed the development of learning-based stereo image compression approaches (Liu, Wang, and Urtasun 2019; Deng et al. 2021; Wödlinger et al. 2022; Zhai et al. 2022; Deng et al. 2023). Given a reference view, these methods, rooted in the predictive coding paradigm, employ disparity-compensated prediction in either pixel or feature space to compress current view. Additionally, recent advancements have introduced a bidirectional coding framework (Lei et al. 2022; Wödlinger et al. 2024) that utilizes a cross-attention mechanism, further enhancing content correlation exploitation between stereo images. Albeit remarkable, the odyssey of stereo image compression research concentrates on producing compact representations through various information flows or network designs before entropy coding. However, an orthogonal direction, which reduces redundancy by devising a superior spatial-disparity entropy model for entropy coding, is rarely explored.
Previous learning-based SIC approaches often adapt existing single image compression solutions, such as hyperprior (Ballé et al. 2018) and auto-regressive prior (Minnen, Ballé, and Toderici 2018), into conditional formats. These adapted entropy models, typically based on convolutional neural network (CNN) architectures, aim to capture spatial-disparity correlations. However, CNNs are inherently limited by their local receptive fields, which hampers their ability to model long-range dependencies. This limitation often leads to a rough estimation of probability distribution and inferior compression performance, especially when the disparity distance between stereo images is large. Consequently, we argue that better coding gains can be achieved as long as we find a more effective spatial-disparity entropy model.
@@
In this paper, we present a stereo image compression framework that centers on a powerful Transformer entropy model to leverage spatial-disparity dependencies between current and previously decoded views. This framework incorporates a straightforward image encoder-decoder pair from He et al. (2022a) to extract latent representations for each view. As illustrated in Fig. 1 (a), we introduce the vanilla masked image modeling (MIM) style (Chang et al. 2022) to construct a content-irrelevant encoder-decoder Transformer entropy model. This model captures spatial information from the context of decoded tokens and disparity information from previous decoded view. Since the bi-directional self-attention in Transformer allows [MASK] tokens to use context from both directions, it supports a non-sequential auto-regressive decoding process and decodes all tokens in the images in a few steps, achieving a good performance-speed trade-off (Chang et al. 2022).
@@
Nevertheless, the vanilla MIM-based Transformer entropy model still exhibits two significant shortcomings. Firstly, it uses uniform [MASK] tokens to indicate the positions of the tokens whose probability distributions are yet to be estimated, which are irrelevant to the image content and uninformative to entropy coding. As a result, it wastes much computation and leads to inferior coding gains. Secondly, it needs an extra Transformer decoder to propagate the prior information via cross-attention. Such a complicated encoder-decoder Transformer architecture only permits an inefficient unidirectional propagation from the prior information to the input tokens of the Transformer decoder.
To overcome these challenges, we develop a novel Content-Aware Masked image modeling Transformer entropy model for Stereo Image Compression (CAMSIC) shown in Fig. 1 (b). Specifically, we introduce content-aware tokens generated by prior information to replace the uninformative [MASK] tokens, forming a new content-aware masked image modeling style, which brings three major benefits. First, the interactions between the proposed content-aware tokens and the estimated tokens provide useful information about each specific position, preventing distraction of previous uninformative [MASK] tokens when updating the token features. Second, it allows a bidirectional information flow between prior information and estimated tokens via self-attention. Compared with static prior information in the vanilla MIM, content-aware tokens, carrying specific prior information, can also be iteratively updated. Thus, it can effectively capture the context information from already estimated tokens and reduce the estimation uncertainty of remaining tokens. Third, since our design naturally embeds prior information propagation into each self-attention operation, we discard the whole Transformer decoder and devise an encoder-only Transformer architecture with low computation overhead. In summary, our main contributions are three-fold:
@@
    We introduce a learning-based stereo image compression framework with a simple image encoder-decoder pair, which uses an elegantly neat but powerful Transformer entropy model based on masked image modeling to exploit the relationship between the left and right images.
@@
    Experimental results show that our proposed method with lower encoding and decoding latency significantly outperforms existing learning-based stereo image compression methods. Detailed ablation studies and analyses validate the effectiveness of each proposed component.
Related Works
@@
Single Image Compression. Traditional standard image codecs, such as JPEG (Wallace 1991), JPEG2000 (Skodras, Christopoulos, and Ebrahimi 2001), BPG (Bellard 2014), and VVC-intra (Bross et al. 2021), typically employ a three-step process involving transformation, quantization, and entropy coding. Regrettably, these steps are optimized independently, which may compromise overall compression performance. In contrast, recent learning-based image compression methods (Ballé, Laparra, and Simoncelli 2017; Ballé et al. 2018; Minnen, Ballé, and Toderici 2018; Cheng et al. 2020; Chen et al. 2021; Zhu, Yang, and Cohen 2021; Chen, Xu, and Wang 2022; Zou, Song, and Zhang 2022; He et al. 2022a; Liu, Sun, and Katto 2023) have demonstrated remarkable success, outperforming traditional methods in terms of rate-distortion (RD) performance. These modern approaches utilize nonlinear transformations (e.g., generalized divisive normalization (Ballé, Laparra, and Simoncelli 2015), attention mechanism (Cheng et al. 2020; Chen et al. 2021), and stacks of residual bottleneck blocks (He et al. 2022a)) to generate concise latent representations. Moreover, a series of advanced entropy models, such as factorized prior (Ballé, Laparra, and Simoncelli 2017), hyperprior (Ballé et al. 2018), and auto-regressive context prior (Minnen, Ballé, and Toderici 2018), are employed to accurately approximate the probability distribution of quantized latent representations. These existing works serve as foundational elements for our stereo image compression solution.
@@
Stereo Image Compression. Traditional multi-view image codecs, such as H.264-based MVC (Vetro, Wiegand, and Sullivan 2011) and H.265-based MV-HEVC (Tech et al. 2015), typically compress stereo images using disparity compensation prediction, which enjoys high compression efficiency but heavily depends on prior knowledge to design hand-crafted modules. Recently, learning-based stereo image compression methods have shown substantial improvements in compression performance. These existing works can be roughly classified into two categories: (1) Unidirectional coding (Liu, Wang, and Urtasun 2019; Deng et al. 2021; Wödlinger et al. 2022; Zhai et al. 2022; Deng et al. 2023) begins with the prediction of a disparity-compensated view in either pixel or feature space, followed by the compression of residuals between the current and predicted views. (2) Bidirectional coding (Lei et al. 2022; Wödlinger et al. 2024; Liu et al. 2025) incorporates a cross-attention mechanism within both the encoder and decoder to exploit the mutual information across stereo images. Apart from the above mentioned joint encoding methods, some recent works (Zhang, Shao, and Zhang 2022; Mital et al. 2023) have ventured into distributed multi-view image coding by using cross-attention alignment to achieve competitive RD performance. Contrary to the prevailing focus on intricate nonlinear transformations by exploring different network structures or information flows, we investigate the application of the Transformer architecture in entropy coding and propose a gracefully simple but powerful Transformer entropy model for stereo image compression. It effectively exploits spatial-disparity correlations, which eliminates the need for complex feature extraction or warping operations, thereby streamlining the compression process and establishing a neat yet potent coding framework.
@@
Masked Image Modeling. Inspired by the masked language modeling (Kenton and Toutanova 2019) in natural language processing, masked image modeling (MIM) has been recently applied to representation learning for vision tasks (He et al. 2022b; Xie et al. 2022; Liu et al. 2022a). It has shown great success on various downstream tasks, including image generation (Chang et al. 2022; Liang et al. 2022), action recognition (Tong et al. 2022; Feichtenhofer et al. 2022), video compression (Xiang, Tian, and Zhang 2023) and video prediction (Gupta et al. 2023). In this paper, we apply the MIM technique to stereo image compression. Note that the vanilla MIM style uses a uniform [MASK] token to take the places of the tokens whose probability distribution are not yet estimated, which wastes much computation on the content-irrelevant [MASK] tokens, thus leading to sub-optimal coding gains. In addition, it resorts to an extra Transformer decoder to propagate the prior information in a unidirectional manner. To address these limitations, we develop a novel content-aware MIM style, which replaces the uninformative [MASK] tokens with content-aware tokens. It not only enables a bidirectional interaction between the prior information and estimated tokens, but also allows us to design an efficient decoder-free Transformer entropy model.
Transformer Architecture. Recent advancements in Transformer architectures have primarily employed two configurations: (i) The encoder-decoder Transformer is tailored for sequence-to-sequence tasks such as machine translation and question answering. Despite its flexibility in generating varying output lengths, it is computationally intensive and requires substantial resources for training due to its dual-module structure. (ii) The decoder-free Transformer is suitable for input understanding tasks (e.g., sentence matching and feature extraction). This structure usually offers quicker training and lower computational demands. Given the effectiveness of Transformer in information fusion and interaction, recent compression practices like VCT (Mentzer et al. 2022) and MIMT (Xiang, Tian, and Zhang 2023) have incorporated encoder-decoder Transformer architectures into entropy models, significantly advancing compression performance. By contrast, we leverage a decoder-free Transformer to achieve a better speed-performance trade-off.
@@
Method
Overview of CAMSIC
Our stereo image compression framework CAMSIC is shown in Fig. 2. Let $\mathcal{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2}\}$ denotes a pair of stereo images, where $\boldsymbol{x}_{v}\in\mathbb{R}^{H\times W\times 3}$ is the image at view $v$ ($v=1,2$) with height $H$ and width $W$. An image encoder $E$ is used to compress each individual view $\boldsymbol{x}_{v}$ to a latent representation $\boldsymbol{y}_{v}\in\mathbb{R}^{h\times w\times d}$, where $h=H/16$, $w=W/16$ and $d$ is the latent dimension. $\boldsymbol{y}_{v}$ is then quantized to $\boldsymbol{\hat{y}}_{v}$ by the quantizer $Q$. Finally, we feed $\boldsymbol{\hat{y}}_{v}$ to an image decoder $D$ to reconstruct the view $\boldsymbol{\hat{x}}_{v}$. Formally, the overall coding procedure can be formulated as:
@@
where $\mathcal{U}(\cdot)$ is uniform noise, * denotes convolution, $\boldsymbol{\mu}_{v}$ and $\boldsymbol{\sigma}_{v}^{2}$ are the estimated means and variances of the Gaussian distributions. When encoding the first view, we expand a learned embedding with the dimension as $\mathbb{R}^{1\times 1\times d}$ to get $\boldsymbol{\hat{y}}_{-1}\in\mathbb{R}^{h\times w\times d}$. In this paper, we refer to prior information as the combination of disparity prior $\boldsymbol{\hat{y}}_{v-1}$ and hyperprior $\boldsymbol{\hat{z}}_{v}$ to distinguish the spatial context from other priors. By leveraging complementary priors, the conditional entropy model can estimate more accurate probability distributions.
Figure 3: Structures of content-irrelevant encoder-decoder Transformer entropy model and our content-aware decoder-free Transformer entropy model. HPD indicates hyperprior decoder. EPM means entropy parameter module. The mask $\boldsymbol{m}_{k}$ starts at zero and is then updated iteratively by a scheduler-based generator during inference. In each iteration, a sinusoidal scheduler computes the number of tokens $n_{k}$ to be encoded/decoded. The mask generator then sets the values of the $n_{k}$ tokens with the lowest estimated bitrates and the previously encoded/decoded tokens as 1 to update $\boldsymbol{m}_{k}$. (a) By using $\boldsymbol{m}_{k}$ to mix the [MASK] token and $\boldsymbol{\hat{y}}_{v}$, a composite context $\boldsymbol{s}_{v}^{k}$ is synthesized and sent to a projection module, where two identity embeddings are added to further different token types. Both the prior information and the composite context are input into a Transformer decoder to estimate the probability distribution parameters $(\boldsymbol{\mu}_{v}^{m_{k}},\boldsymbol{\sigma}_{v}^{m_{k}})$ for the masked tokens. (b) The [MASK] token is replaced by the context-aware tokens $\boldsymbol{c}_{v}$ to form the composite context $\boldsymbol{u}_{v}^{k}$. This context is then input into a projection module with two identity embeddings and a Transformer encoder to output the estimated Gaussian distribution parameters $(\boldsymbol{\mu}_{v}^{m_{k}},\boldsymbol{\sigma}_{v}^{m_{k}})$.
Content-irrelevant Encoder-decoder Transformer Entropy Model
@@
Encoder-decoder Transformer. Since these [MASK] tokens are irrelevant to the image content, a Transformer decoder is required to propagate the content-related prior information to the [MASK] tokens through cross-attention. Fig. 3 (a) shows the architecture of an encoder-decoder Transformer entropy model based on content-irrelevant MIM. Specifically, we use a prior fusion network and a Transformer encoder to integrate the hyperprior $\boldsymbol{\hat{z}}_{v}$ and disparity prior $\boldsymbol{\hat{y}}_{v-1}$ for producing prior features. During training, we randomly generate a Boolean mask to composite the spatial context $\boldsymbol{s}_{v}$. The prior features and unmasked spatial tokens convey their information to the [MASK] tokens via cross-attention and self-attention in the Transformer decoder, respectively. Finally, we feed the feature $\boldsymbol{f}_{v}^{m}$ produced by the Transformer decoder into an entropy parameter module to estimate the probability distribution parameters $(\boldsymbol{\mu}_{v}^{m},\boldsymbol{\sigma}_{v}^{m})$.
Content-aware Decoder-free Transformer Entropy Model
Content-aware Masked Image Modeling. The vanilla MIM uses a content-irrelevant [MASK] token to occupy the places of the tokens whose probability distribution are not yet estimated, which is not the most efficient and effective choice for entropy coding. On the one hand, interacting with the uninformative [MASK] tokens may distract the specific features of the estimated tokens during the repetitive self-attention operations. On the other hand, the encoder-decoder Transformer architecture is complicated but only allows a unidirectional propagation from prior information to the input tokens of the Transformer decoder. The prior information is fixed once extracted and cannot utilize the updated spatial context information from the already estimated tokens. Consequently, it is difficult for the entropy model to make full use of the power of Transformer architecture.
@@
Decoder-free Transformer. Since the prior information has sufficiently interacted with the estimated tokens in the Transformer encoder, an explicit Transformer decoder is no longer needed. Thus, we discard the whole Transformer decoder and build an encoder-only Transformer architecture. As shown in Fig. 3 (b), it comprises two parts, i.e., a prior generation network that produces content-aware tokens $\boldsymbol{c}_{v}$ based on different priors and a decoder-free Transformer that estimates the probability distributions of tokens. Note that we enable relative position embedding (Liu et al. 2021) in the Transformer and assign two identity embeddings to the token sequence to further distinguish different types of tokens, which effectively promote the token interactions.
Training and Inference
@@
During inference, the probability distributions of the tokens are iteratively estimated following a sinusoidal decoded ratio scheduler (Chang et al. 2022). Let $R(\hat{\boldsymbol{y}}_{v}^{i,k})$ denotes the estimated bitrate of $i$-th token in $\hat{\boldsymbol{y}}_{v}$ during $k$-th iteration. As shown in Fig. 3, the mask $\boldsymbol{m}_{0}$ is initially zero during inference. In each iteration, we use the composite context $\boldsymbol{u}_{v}^{k}$ to estimate the bitrate $R(\hat{\boldsymbol{y}}_{v}^{i,k})$. The mask generator employs a scheduling function to calculate the number of tokens $n_{k}$ to be encoded/decoded, and selects those $n_{k}$ tokens that consume the fewest bits to update the mask $\boldsymbol{m}_{k+1}$, setting the values in the $n_{k}$ tokens and previous encoded/decoded tokens to 1. Finally, $\boldsymbol{u}_{v}^{k+1}$ integrates $\hat{\boldsymbol{y}}_{v}$ and $\boldsymbol{c}_{v}$ using $\boldsymbol{m}_{k+1}$ for next iteration. This iterative process ensures consistency between encoder and decoder, producing identical $(\boldsymbol{\mu}^{m_{k}}_{v},\boldsymbol{\sigma}^{m_{k}}_{v})$. More details about iterative decoding process are available in the appendix.
Quantization. During training, since the quantization operation is not differentiable, we leverage a mixed quantizer to allow end-to-end optimization. To be specific, it adds $\boldsymbol{y}_{v}$ and a uniform noise $\mathcal{U}(-\frac{1}{2},\frac{1}{2})$ together as the input to the entropy model, but rounds $\boldsymbol{y}_{v}$ as the input to the decoder based on straight-through estimation (Bengio, Léonard, and Courville 2013). We refer readers to Minnen and Singh (2020) for more details.